1. Given m,,m,, m,0 and H, (same for m, and m,) find the equation of motion ofthe system sho wn: T1 T m 25.0° Using m, = 10.0kg , m2 = 5.00kg , m, = 3.00kg , θ = 25.0, and both ropes. = 0.200 find the tension in Find the tension in the two cables I'm confused on how to draw the free body diagrams
Answer #1

equation for mass m1 is

T1 = m1g = 9.8 *10 = 9.8 N

equation for mass m2 is

T2 + uk*m2*g = T1

from this we get T2 = 88.2 N

equation for mass m3 is

T2 + uk*m3*g*cos25 = m3*g*sin25

there is some error in image uploading, so its difficult to share free body diagram